OpenStudy (anonymous):

Find the major axis for the ellipse with the given equation. x^2 + 16y^2 - 96y + 128 = 0 A)8 B)16

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

My teacher said it wasn't 4 though

OpenStudy (anonymous):

Then I don't know... I was using mathway..

OpenStudy (anonymous):

Oh ok thanks anyway

OpenStudy (irishboy123):

try taking it this way: \(x^2 + 16y^2 - 96y + 128 = 0\) \(x^2 + 16(y^2 - 6y) + 128 = 0\) \(x^2 + 16[(y- 3)^2 -9] + 128 = 0\) etc etc ......until you have it in the standard form, which here is: \( \displaystyle{\left(\frac{x}{a}\right)^2 + \left(\frac{y - y_o}{b}\right)^2} = 1\) there's no point asking computers to do it for you :p